### v1

@@ Abstract
Trade restrictions, the COVID-19 pandemic, and geopolitical conflicts has significantly exposed vulnerabilities within traditional global supply chains. These events underscore the need for organisations to establish more resilient and flexible supply chains. To address these challenges, the concept of the autonomous supply chain (ASC), characterised by predictive and self-decision-making capabilities, has recently emerged as promising solution. However, research on ASCs is relatively limited, with no existing studies on their implementations. This paper aims to address this gap by presenting an implementation of ASC using a multi-agent approach. It proposes a methodology for the analysis and design of such an agent-based ASC system (A2SC). This paper provides a concrete case study, the autonomous meat supply chain, which showcases the practical implementation of the A2SC system using the proposed methodology. Additionally, a system architecture and a toolkit for developing A2SC systems are presented. Despite with limitations, this paper demonstrates a promising approach for implementing an effective ASC system.
@@ 1 Introduction
Trade restrictions, the COVID-19 pandemic, and geopolitical conflicts has seriously exposed vulnerabilities within traditional global supply chains [Handfield et al., 2020, Shih, 2020], highlighting the need for organisations to establish more resilient and flexible supply chains. Indeed, traditional supply chains exhibit limited agility and resilience, with over 70% of surveyed supply chain leaders acknowledging the need for enhanced agility and operational resilience [NelsonHall, 2021]. The model of autonomous supply chains (ASC) has been proposed as a concept solution to achieve supply chain resilience and structural flexibility [Xu et al., 2022].
@@
Traditional supply chains still heavily rely on manual processes, which can often introduce communication delays and human errors, ultimately leading to increased costs [NelsonHall, 2021]. Intelligent digital technologies, such as AI and robotics, are penetrating into supply chain domain, automating various processes, particularly those error-prone and routine tasks. However, these automation efforts mainly focuses on individual processes, with limited integration between processes, let alone inter-firm integration. The needs for better supply chain performance require more horizontal and vertical integration [Frohlich and Westbrook, 2001, Stevens and Johnson, 2016]. ASCs emphases interconnection, integration, automation, and self-decision-making. Within this framework, the supply chain has capacity to predict disruptions and automatically adapt to changes through reconfiguration and adjustments. Consequently, ASCs offer a promising solution to address challenges in an uncertain and turbulent business environment.
@@
The concept of supply chains equipped with such predictive and self-adaptive capabilities have been a topic of discussion for many years. A “smarter” supply chain, capable of autonomous learning and decision-making without human intervention, was put forth by Butner [2010]. This envisioned supply chain was characterised as instrumented, interconnected, and intelligent. Furthermore, Wu et al. [2016] conducted a literature review on the impact of digital technologies on SCM and proposed three additional defining characteristics for this emerging supply chain: automation, integration, and innovation. Expanding on these ideas, Calatayud et al. [2019] proposed the concept of the self-thinking supply chain, which is distinguished by autonomous and predictive capabilities relying on IoT and AI-enabled connectivity. Moreover, Xu et al. [2022] extended these concepts by formally proposing the autonomous supply chain model. This work extensively delved into the theoretical aspects of ASCs, including their formal definition, defining characteristics, and a set of auxiliary concepts. Instead of solely outlining the characteristics of this intelligent supply chain, it presented a five-layer conceptual model and a seven-level supply chain autonomy reference model. These models serve as practical guidelines for the implementation of ASC systems.
However, despite the existence of initial concepts and theoretical studies in literature, the development of ASCs remains in its nascent stages. Notably, none of these works have extensively addressed the technical aspects of ASCs. This paper, therefore, aims to bridge this gap by focusing on the technical implementation of ASCs. While Xu et al. [2022] presented a short high-level overview of an ASC implementation case, this paper delves comprehensively into all facets of implementing an ASC system, employing a Multi-Agent System (MAS) approach.
@@
A supply chain consists of a network of business entities engaged in various activities that transform raw materials into finished products and distribute them to end customers [Christopher, 2016]. These activities include demand forecast, material procurement, inventory management, and order fulfilment. Effectively managing the flow of goods and its associated information and financial flows is crucial in supply chain management (SCM). To achieve this automatically, interoperability between independent and geographically distributed entities must be facilitated. MAS approaches, which achieved distributed problem-solving through multiple intelligent agents interacting within a shared environment, are inherently well-suited for architecturally modelling supply chains [Wooldridge and Jennings, 1995, Fox et al., 2001, Xu et al., 2023]. These agents can be seen as an independent, goal-oriented software process that operates autonomously, communicating and coordinating with other agents as needed. This paper investigates the creation of an agent-based ASC (A2SC) system that facilitates interoperability and automation among distributed heterogeneous supply chain entities.
As described in Xu et al. [2022], entirely automating a supply chain involves replacing manual processes by automated ones and establishing the connections between these automated processes. These automated processes, which are constituent building blocks for ASCs, can be managed through a set of structural entities defined in Xu et al. [2022]. Moreover, these structural entities are responsible for managing the three flows that traverse these interconnected processes. In A2SCs, these structural entities can be represented by software agents, autonomously managing its internal entities and interacting with external entities.
Despite agent-based technology being proposed for use in SCM for over two decades, it current technical readiness level (TRL) remains relatively low, and its industrial adoption is limited [Xu et al., 2021]. This paper aims to investigate fundamental issues relevant to the design and development of a practical A2SC system, including its analysis, design, and implementation.
The main contributions of this paper can be summarised as follows:
  1. (1)
@@
    We investigate the application of the MAS approach in supply chain domain and introduce an Agent-based Autonomous Supply Chain (A2SC).
  2. (2)
@@
    We present a methodology designed for the analysis and design of A2SC systems, and detail the realisation of A2Sc systems using this methodology.
  3. (3)
@@
    We present a detailed concrete case study, the autonomous meat supply chain, demonstrating the practical implementation of an A2SC system.
  4. (4)
@@
    We report a system architecture and a toolkit for A2SC prototype development, potential offering guidance for similar decentralised systems.
@@
The rest of this paper is organised as follows. Section 2 reviews related work. Section 3 presents agent-based autonomous supply chains and a methodology for its analysis and design. Section 4 details an A2SC implementation case study. Section 5 discusses the limitation and implications of this work. Finally, Sections 6 concludes this paper and discusses future work.
@@ 2 Related Work
Given that ASCs remain at its early stages, tehre are limited existing literature related to their technical implementation. This section reviews the work related to the use of MAS approaches in SCM, which is relevant for implementing A2SCs.
The adoption of MAS approaches in SCM can be traced back to the development of the Integrated Supply Chain Management System (ISCM) at the University of Toronto in the early 1990s [Fox et al., 1993]. The ISCM was composed of a collection of software agents, with each responsible for one or more activities within the supply chain. These agents coordinated with one another to plan and execute supply chain functions. Fox et al. [2001] investigated key issues and provided detailed solutions for constructing such an agent-based supply chain architecture. Similarly, Swaminathan et al. [1998], another early representative work, employed a multi-agent framework to model the dynamic aspects of supply chains. This study created a library of software agents that capture generic supply chain processes and concepts, providing a modular and reusable framework for developing a wide range of realistic supply chain models. In contrast to Fox et al. [1993, 2001], which decomposed agents based on supply chain functions, Swaminathan et al. [1998] defined agents by considering the structural elements in the supply chain. The proposed framework in Swaminathan et al. [1998] allows models to address issues at both tactical (coordination) and strategic (configuration) levels. These early pioneering works laid the foundation for the subsequent development of multi-agent frameworks to address SCM challenges.
@@
Since then, numerous applications of agent technology in SCM have emerged, which can be broadly categorised into four groups. The first category involves agent frameworks designed for generic or specific supply chains. These efforts aim to identify common characteristics of generic supply chains or individual industry-specific supply chain, providing a framework for constructing MAS for supply chain operations. For example, Sadeh et al. [2001] introduced a MASCOT architecture based on the blackboard architectural model [Erman et al., 1980]. This architecture providing an generic and reconfigurable agent framework for supply chain coordination and dynamic formation. Julka et al. [2002] defined a unified framework for modelling entire supply chain network in an object-oriented manner, with supply chain entities represented as agents and materials and information represented as objects. [Lou et al., 2004] proposed an agent architecture that consists of mediator and functional agents for agile SCM. Agent frameworks have also been applied in specific supply chains, such as fashion [Lo et al., 2008], wine [Kumar et al., 2013], outsourcing SME manufacturing [Kumar et al., 2013], and construction [Xue et al., 2005].
@@
The second category focuses on utilising agent-based approaches to address dynamic supply chain formation and configuration. For example, Walsh and Wellman [1999] discussed the issues and models relevant to automated supply chain formation using an agent-based approach. Chen et al. [1999] introduced a MAS that utilises a set of performatives for pair-wise negotiation, enabling dynamic supply chain generation in an open environment through automated or semi-automated negotiation. Similar to Fox et al. [1993], this MAS consists of two types of agents: functional agents and information agents. Kim and Cho [2010] investigated the dynamic supply chain formation problem based on the Single Machine Earliness/Tardiness (SET) model and employed agent negotiation as a means to address this problem in a three-tier supply chain setting. Akanle and Zhang [2008] developed an agent-based model to manage customer demand over a period of time by optimising supply chain configuration. Meanwhile, Ameri and McArthur [2013] presented an agent-based framework, the Digital Manufacturing Market (DMM), for handling supply chain configuration. The DMM combined semantic web technology and agent technology to enable automated service discovery and matching, thereby facilitating dynamic configuration of a two-tier supply chain that only involves supplier and customers.
@@
The third category relates to coordination and negotiation in SCM. While supply chain formation and configuration relate to the structural aspects of a supply chain, coordination and negotiation address its dynamical aspects [Swaminathan and Tayur, 2003]. Effective coordination among supply chain functions is essential for efficient operations. Barbuceanu and Fox [1996] tackled the coordination problem by explicitly representing interaction knowledge among agents and introduced COOL, a structured conversion-based communication language for supply chain coordination. Sadeh et al. [2001] developed the MASCOT architecture, which offers protocols for both horizontal and vertical coordination in supply chains. Xue et al. [2005] proposed an agent framework for construction supply chain coordination, based on multi-issue negotiation using multi-attribute utility theory (MAUT). Wong and Fang [2010] presented a contract-net-like multilateral and multi-issue protocol designed for handling buyers-seller bargaining and interaction negotiation in SCM. This protocol also adopts the MAUT approach to model utility functions during negotiations. Various general negotiation approaches suitable for SCM have been proposed, including the approach by Chen and Weiss [2015].
@@
While the previous three categories focus on constructing agent-based SCM systems, the last category, which is main focus of most literature, aims to enhance specific functions or processes through MAS approaches, such as demand forecast [Liang and Huang, 2006], cost management [Fu and Fu, 2015], supplier selection [Ghadimi et al., 2018, Yu and Wong, 2015], scheduling [Aminzadegan et al., 2019], inventory planning and control [Ito and Mousavi Jahan Abadi, 2002, Kazemi et al., 2009, Kim et al., 2010], and risk and event management [Ahn et al., 2003, Bodendorf and Zimmermann, 2005, Giannakis and Louis, 2011, Bearzotti et al., 2012].
While there is a substantial body of literature on agent-based SCM (for a comprehensive review, see Xu et al. [2021] ), none of them focus on achieving the automation across the entire supply chain, including in both decision making and execution aspects. We in this paper aim to address this gap, developing methodology to design ASC systems using MAS approaches and presenting a technical framework for implementing a concrete A2SC prototype system.
3 Agent-based Autonomous Supply Chain (A2SC)
@@
The adoption of technologies in the Fourth Industrial Revolution (Industry 4.0) [World Economic Forum, 2017] for transforming supply chains has gained significant attention [BlueYonder, 2020, Calatayud et al., 2019], with existing academic and industrial efforts mainly focusing on the application of digital technologies like AI and advanced robotics to specific functions, such as demand forecasting and pallet picking, rather than fully automating entire processes.
@@
Automated functions are building blocks of fully automated processes, which, in turn, collectively form an entirely automated system. Within a supply chain, numerous functions require automation. Automating these specific functions needs multidisciplinary efforts and technological advancements, which is beyond of the scope of this paper. Instead, this paper focuses on integrating these “already-automated” functions and facilitating entire processes. Effective coordination mechanisms, both within individual business entities and across them, are crucial for achieving a seamless ASC. In this section, we describe MAS approaches, with a particular focus on the methodology for analysing and designing A2SC systems.

### v2

@@ Abstract
Trade restrictions, the COVID-19 pandemic, and geopolitical conflicts have significantly exposed vulnerabilities within traditional global supply chains. These events underscore the need for organisations to establish more resilient and flexible supply chains. To address these challenges, the concept of the autonomous supply chain (ASC), characterised by predictive and self-decision-making capabilities, has recently emerged as a promising solution. However, research on ASCs is relatively limited, with no existing studies specifically focusing on their implementations. This paper aims to address this gap by presenting an implementation of ASC using a multi-agent approach. It presents a methodology for the analysis and design of such an agent-based ASC system (A2SC). This paper provides a concrete case study, the autonomous meat supply chain, which showcases the practical implementation of the A2SC system using the proposed methodology. Additionally, a system architecture and a toolkit for developing such A2SC systems are presented. Despite limitations, this work demonstrates a promising approach for implementing an effective ASC system.
@@ 1 Introduction
Trade restrictions, the COVID-19 pandemic, and geopolitical conflicts have seriously exposed vulnerabilities within traditional global supply chains (Handfield et al., 2020; Shih, 2020), highlighting the need for organisations to establish more resilient and flexible supply chains. Indeed, traditional supply chains exhibit limited agility and resilience, with over 70% of surveyed supply chain leaders acknowledging the need for enhanced agility and operational resilience (NelsonHall, 2021). The model of autonomous supply chains (ASC) has been proposed as a concept solution to achieve supply chain resilience and structural flexibility (Xu et al., 2023b).
@@
Traditional supply chains still heavily rely on manual processes, which can often introduce communication delays and human errors, ultimately leading to increased costs (NelsonHall, 2021). Intelligent digital technologies, such as AI and robotics, are penetrating into the supply chain domain, automating various processes, particularly those error-prone and routine tasks. However, these automation efforts mainly focus on individual processes, with limited integration between processes, let alone inter-firm integration. The needs for better supply chain performance require more horizontal and vertical integration (Frohlich and Westbrook, 2001; Stevens and Johnson, 2016). ASCs emphasise interconnection, integration, automation, and self-decision-making. Within this framework, the supply chain has capacity to predict disruptions and automatically adapt to changes through reconfiguration and adjustments. Consequently, ASCs offer a promising solution to address challenges in an uncertain and turbulent business environment.
@@
The concept of supply chains equipped with such predictive and self-adaptive capabilities have been a topic of discussion for many years. A “smarter” supply chain, capable of autonomous learning and decision-making without human intervention, was proposed by Butner (2010). This envisioned supply chain was characterised as instrumented, interconnected, and intelligent. Wu et al. (2016) conducted a literature review on the impact of digital technologies on supply chain management (SCM) and proposed three additional defining characteristics for this emerging supply chain: automation, integration, and innovation. Expanding on these ideas, Calatayud et al. (2019) proposed the concept of the self-thinking supply chain, which is distinguished by autonomous and predictive capabilities relying on IoT and AI-enabled connectivity. Moreover, Xu et al. (2023b) extended these concepts by formally proposing the autonomous supply chain model. This work extensively delved into the theoretical aspects of ASCs, including their formal definition, defining characteristics, and a set of auxiliary concepts. Instead of solely outlining the characteristics of this intelligent supply chain, it presented a five-layer conceptual model and a seven-level supply chain autonomy reference model. These conceptual models can be used for providing guidelines for implementing ASC systems.
@@
However, despite the existence of initial concepts and theoretical studies in the literature, the development of ASCs remains in its early stages. A supply chain consists of a network of business entities engaged in various activities that transform raw materials into finished products and distribute them to customers (Christopher, 2016). These activities include demand forecast, supplier selection, material procurement, inventory management, and order fulfilment. Effectively managing the flow of goods and its associated information and financial flows is crucial for SCM. To achieve autonomous SCM, interoperability between independent and geographically distributed entities must be facilitated. The multi-agent system (MAS) approach, which achieves distributed problem-solving through multiple intelligent agents interacting within a shared environment, are inherently well-suited for architecturally modelling supply chains (Wooldridge and Jennings, 1995; Fox et al., 2001; Xu et al., 2023c). While the MAS approach has been long recognised by academics and practitioners for its potential to automate and autonomise decision-making in SCM, its applications remain limited mainly to addressing the automation of specific functions or processes (Xu et al., 2021; Nitsche et al., 2023).
For instance, Ameri and McArthur (2013) introduced an agent-based digital manufacturing marketplace, enabling service consumers to autonomously search for and connect with corresponding service providers based on a predefined similarity metric. Despite facilitating the creation of configurable supply chains, this service matchmaking still targets the automation of a single function — supplier selection. In a more recent study, Nitsche et al. (2023) conducted interviews with a cohort of supply chain experts, identifying eleven MAS implementation use cases in SCM, all primarily centred around automating specific functions rather than the automation of physical flows and their associated information flows. Furthermore, these use cases heavily rely on agent simulation/development platforms such as JADE (Bellifemine et al., 2005) for testing and evaluation, which deviates from realistic settings and exhibits limited compatibility with legacy systems (Xu et al., 2021). Consequently, this reliance contributes to the slow industrial adoption of MAS for automation and autonomisation in SCM.
Additionally, current research rarely focuses on automating the entire SCM, which involves managing the three essential flows from material procurement to the final delivery of finished product to customers. Autonomous management of these three flows is crucial for achieving ASCs (Xu et al., 2023b). This paper, therefore, aims to bridge this gap by systematically exploring the development of ASCs with integrated processes. While Xu et al. (2023b) presented a brief overview of an ASC implementation case, this paper delves comprehensively into all facets of implementing an ASC system, employing a multi-agent system (MAS) approach. Compared to existing studies, this work distinguishes itself by 1) adopting a software engineering methodology — Gaia (Wooldridge et al., 2000) and its variant ROADMAP (Juan et al., 2002) — as system design and analysis approaches; 2) tackling the automation of the entire process from sourcing to delivery to customers; and 3) implementing a software system prototype rather than a simulation. Specifically, this work investigates the development of an agent-based ASC system, called A2SC, that facilitates interoperability and automation among distributed supply chain entities.
In this paper, we scrutinise the fundamental issues relevant to the design and development of a practical A2SC system, including its analysis, design, and implementation. The main contributions of this paper can be summarised as follows:
@@
  • We investigate the application of the MAS approach in automating integrated processes and introduce an agent-based ASC.
@@
  • We present a methodology designed for A2SC analysis and design, and detail the realisation of A2SC systems using this methodology.
@@
  • We detail a concrete case study — the autonomous meat supply chain — demonstrating the technical implementation of an A2SC system.
@@
  • We report a system architecture and a toolkit for A2SC prototype development, providing guidance for implementing similar decentralised systems.
@@
The rest of this paper is organised as follows. Section 2 reviews related work. Section 3 presents A2SCs and a methodology for their analysis and design. Section 4 details an A2SC implementation case study. Section 5 discusses the limitation and implications of this work. Finally, Section 6 concludes this paper and discusses future work.
@@ 2 Related Work
While ASCs are still in their early stage, there exists a body of research pertinent to their development. We classify these work into two categories: conceptual discussions focusing on the transition towards ASCs and related technical attempts aiming at automating and autonomising SCM through MAS. In this section, we provide a review of these two categories of related work.
2.1 Transition to ASCs
The initial discussion of ASCs can be traced back the concept of the “smarter supply chain” proposed by Butner (2010) — a supply chain capable of autonomous learning and decision-making without human intervention. Despite the description of the characteristics of such supply chains in this article, their conceptual or theoretical aspects of ASCs have not been extensively studied. Wu et al. (2016) further developed the concept and investigated various technologies relevant to achieving a “smart supply chain”. Built upon both physical and information connectivity (Calatayud and Katz, 2017), Calatayud et al. (2019) proposed the “self-think supply chain”, characterised by autonomous and predictive capabilities. Similar to Wu et al. (2016), this work defines the characteristics of this type of supply chains and identifies key enabling technologies to achieve them. Nitsche (2021); Nitsche et al. (2021) proposed a conceptual framework that specifies application areas and antecedents for realising automation in SCM and logistics. Compared to previous studies, this work concretises concepts and centres around conceptualisation of the automation of physical and information processes. Although this conceptual framework provides a common basis for further discussion on automation between research and practice, it falls short of guiding the technical development of ASCs.
Xu et al. (2023b) delved deeper into the conceptualisation of ASCs, proposing a set of conceptual artefacts, including structural entities and the MIISI model, to guide ASC implementation. As described in Nitsche (2021); Xu et al. (2023b), fully automating a supply chain involves replacing manual processes by automated ones and establishing connections between these automated processes. These automated processes served as the building blocks for ASCs and can be managed through structural entities. These structural entities are entities along a supply chain network that assemble and control the flows of materials, information, and finance. They are responsible for managing the three flows that traverse these interconnected processes. In A2SCs, these structural entities can be represented by software agents that autonomously manage their internal units and interacting with external entities. All these works focus on conceptual aspects of ASCs; however, none of them directly deal with their technical implementation. We review the work related to their technical implementation in the next section.
2.2 Realisation of ASCs
A supply chain process requires numerous decisions among a multitude of actors across a supply chain. The MAS approach, which involves a collection of software agents interacting with each other to solve a complex problem, is therefore well-suited to facilitate the automation and autonomisation of these processes (Dorri et al., 2018; Xu et al., 2021; Nitsche et al., 2023). Since its initial adoption in SCM (Fox et al., 1993), yet still under-researched (Dorri et al., 2018), this approach has been widely explored in this domain.
@@ 2.2 Realisation of ASCs
These applications can be broadly grouped into four categories: agent frameworks, dynamic supply chain formation and configuration, coordination and negotiation, and enhancement of specific supply chain functions. The first category focuses on identifying common characteristics of generic or specific supply chains, providing an architectural framework for constructing agent-based SCM. For instance, Sadeh et al. (2001) introduced a MASCOT architecture based on the blackboard architecture (Erman et al., 1980). Julka et al. (2002) defined a unified framework for modelling the entire supply chain network in an object-oriented manner, representing supply chain entities as agents and materials and information represented as objects. Lou et al. (2004) proposed an agent architecture consisting of mediator and functional agents for agile SCM. Additionally, agent frameworks designed for supply chains of specific industries, such as fashion (Lo et al., 2008), wine (Kumar et al., 2013), and outsourcing SME manufacturing (Kumar et al., 2013), have also been proposed.
@@
The second category utilises the MAS approach to address dynamic supply chain formation and configuration. Chen et al. (1999) introduced an MAS that employs a set of performatives for pair-wise negotiation, facilitating dynamic supply chain generation in an open environment through automated or semi-automated negotiation. Akanle and Zhang (2008) developed an agent-based model to manage customer demand by optimising supply chain configuration. Kim and Cho (2010) investigated the dynamic supply chain formation problem based on the single machine earliness/tardiness model and utilised agent negotiation to address this issue in a three-tier supply chain setting. Meanwhile, Ameri and McArthur (2013) presented a digital manufacturing market for matching service providers and consumers, enabling semantic matchmaking and thereby facilitating supply chain reconfiguration.
@@
The third category specifically addresses coordination and negotiation within SCM. While the second category focuses on the structural aspects of supply chains, this category deals with their dynamical aspects (Swaminathan and Tayur, 2003) — coordination and negotiation. Barbuceanu and Fox (1996) addressed the coordination problem by explicitly representing interaction knowledge among agents and introduced COOL, a structured conversion-based communication language for supply chain coordination. Sadeh et al. (2001) developed the MASCOT architecture, which provides protocols for both horizontal and vertical coordination in supply chains. Xue et al. (2005) proposed an agent framework for coordinating construction supply chains, leveraging multi-issue negotiation using multi-attribute utility theory (MAUT). Wong and Fang (2010) presented a contract-net-like multilateral and multi-issue protocol designed for handling buyers-seller bargaining and interaction negotiation in SCM. This protocol also employs the MAUT approach to model utility functions during negotiations. Additionally, various general negotiation approaches suitable for SCM have been proposed, including the approach by Chen and Weiss (2015).
@@
The last category, which is main focus of most literature, aims to improve specific supply chain functions or processes through the MAS approach. These include enhancing demand forecast (Liang and Huang, 2006), managing costs (Fu and Fu, 2015), selecting suppliers (Ghadimi et al., 2018; Ameri and McArthur, 2013; Yu and Wong, 2015), scheduling (Aminzadegan et al., 2019), planning and controlling inventory (Ito and Mousavi Jahan Abadi, 2002; Kazemi et al., 2009; Kim et al., 2010), and managing risks and events (Ahn et al., 2003; Bodendorf and Zimmermann, 2005; Giannakis and Louis, 2011; Bearzotti et al., 2012; Proselkov et al., 2024).
As discussed above, existing research has mainly focused on using the MAS approach to address specific functions or processes within supply chains, with less attention given to automating integrated processes. However, achieving ASCs would require seamlessly integrating multiple processes across company boundaries to autonomously managing the three supply chain flows (Xu et al., 2023b; Nitsche et al., 2023). Additionally, existing work has relatively low technical readiness (Karnouskos and Leitao, 2016) and the implementation of more complex problems are rare (Nitsche et al., 2023). This paper thus aims to address these gaps. We study on a more complex scenario involving integrated processes and multiple tiers of supply chain stakeholders, automating its entire physical flows and their associated information flows using the MAS approach. Moreover, we adopt a software engineering methodology for system design and analysis, aiming to achieve higher technical readiness and reduce the divide between academic exploration and industrial adoption.
3 Agent-based Autonomous Supply Chain
@@
The adoption of the Industry 4.0 technologies (World Economic Forum, 2017) for transforming supply chains has gained significant attention (BlueYonder, 2020; Calatayud et al., 2019), with existing academic and industrial efforts mainly focusing on using AI, IoT, and advanced robotics for enhancing specific functions, such as demand forecasting and pallet picking, rather than automating integrated processes.
@@
Automated functions are “building blocks” of an integrated automated process, which, in turn, collectively form an entire automated system. An ASC involves numerous automated processes, autonomously managing the three flows. Automating all specific functions needs both technological and managerial efforts (Nitsche et al., 2023), which is beyond of the scope of this paper. Therefore, this paper focuses on integrating these “already-automated” functions and facilitating entire processes. Effective coordination mechanisms, both within individual business entities and across them, are crucial for achieving a seamless ASC. This section describes the MAS approach, with a particular focus on the methodology for analysing and designing A2SC systems.
3.1 Agents and Multi-Agent System